OpenStudy (anonymous):

Use Newton's method to get an approximation to sqrt(12) that is accurate to four decimal places.

OpenStudy (anonymous):

newton's method in this case is the same as "mechanic's rule'. pick a number close to \[\sqrt{12}\] say 3 \[x_1=3\] \[x_2=\frac{1}{2}(3+\frac{12}{3})\]

OpenStudy (anonymous):

then repeat the process. in \[x_{n+1}=\frac{1}{2}(x_n+\frac{12}{x_n})\]

OpenStudy (anonymous):

once you get the same decimal twice to 4 places stop. incidentally this is pretty much how your calculator does it.

myininaya (myininaya):

\[x=\sqrt{12}\] then \[x^2-12=0\] let \[f(x)=x^2-12\] so \[f'(x)=2x\] we choose a x=3 like satellite did because we know this already a close apporximation to \[\sqrt{12}\] so do \[x_{n+1}=x_n-\frac{f(n)}{f'(n)}\]

myininaya (myininaya):

\[x_1=3-\frac{f(3)}{f'(3)}=3-(-3)/(6)=3+1/2=3.5\]

OpenStudy (anonymous):

idea is this. you pick a number \[x_1\]close to \[\sqrt{a}\] it is either too big or two small then consider \[\frac{a}{x_1}\] which is also close since \[x_1\times \frac{a}{x_1}=a\] but on the other side of \[\sqrt{a}\] meaning if \[x_1\] is too big then \[\frac{a}{x_1}\] is too small and vice versa. then just take the average of the two numbers via \[x_2=\frac{1}{2}(x_1+\frac{a}{x_1})\] and repeat the process until you are satisfied with the resulg

myininaya (myininaya):

\[x_2=3.5-\frac{f(3.5)}{f'(3.5)}=3.5-\frac{.25}{7}=3.464285714\] \[x_3=3.464285714-\frac{f(3.464285714)}{f'(3.464285714)}=3.46410162\] \[x_4=3.46410162-\frac{f(3.46410162)}{f'(3.46410162)}=3.464101615\]

OpenStudy (anonymous):

to see that they are the same, write what myininnay wrote \[x_{n+1}=x_n-\frac{f(x_n)}{f'(x_n)}\] \[x_{n+1}=x_n-\frac{x_n^2-12}{2\times x_n}\] \[x_{n+1}=\frac{2x_n^2-x_n^2+12}{2x_n^2}\] \[x_{n+1}=\frac{x_n^2+12}{2x_n}=\frac{1}{2}(x_n+\frac{12}{x_n})\]

myininaya (myininaya):

look how close my number is to the sqrt{12} isn't it a beauty :)

OpenStudy (anonymous):

this subscript business is a pain hello myininaya !

myininaya (myininaya):

hey

OpenStudy (anonymous):

yes it is "error squaring'

OpenStudy (anonymous):

meaning if you have 2 place accuracy on step 3 then you are guaranteed at least 4 place accuracy on step 4 and 8 place on step 5 etc

myininaya (myininaya):

i know you said only 4 decimals but why not the best approximation i can give with the calculator that i have

OpenStudy (anonymous):

\[\color{blue}{\text{did you stay up all night?}}\]

myininaya (myininaya):

yes i did

OpenStudy (anonymous):

it is remarkable how fast you converge using newton's method aka mechanics rule. this is ancient btw

myininaya (myininaya):

but you could have stopped at x_2 or x_3

OpenStudy (anonymous):

way before calculus

myininaya (myininaya):

i don't think linearization is as awesome what do you think?

OpenStudy (anonymous):

in fact is it almost common sense. pick a number, find the corresponding one so that xy = a and then take the average

OpenStudy (anonymous):

i like this because it required just thinking. i want root 12 i pick 3, which is too small

OpenStudy (anonymous):

then i say ok well 3*4=12 so 3 is too small, 4 is too big, take the average and get 3.5

myininaya (myininaya):

well we knew it was to small because 3^2=9 which is smaller than 12

OpenStudy (anonymous):

that is all this does at each step.

OpenStudy (anonymous):

now i have 3.5 then i say ok 12/3.5 3.4285..

myininaya (myininaya):

lol mnad has something to say

OpenStudy (anonymous):

3.5 is too big, 3.4385 is too small, take the average and get ...

OpenStudy (anonymous):

thank you for the elaborate answers, guys. Both are great, I went with satellite's for simplicity reasons. If I provide any more than 4 decimals then I risk having points taken off :/ My professor is a bit of a harsh grader

myininaya (myininaya):

lol ok

OpenStudy (anonymous):

well if you have time convince yourself that they are the same. and when you are bored do this to find \[\sqrt{2}\]

OpenStudy (anonymous):

you will see some patterns in your answer if you do it with pencil and paper, patterns that get lost in decimals

myininaya (myininaya):

have fun!

OpenStudy (anonymous):

It will be extremely useful for my final exam review tomorrow, I'm sure.

myininaya (myininaya):

also newton method doesn't always succeed

OpenStudy (anonymous):

in other words use fractions instead of decimals. for root two use \[x_{n+1}=\frac{1}{2}(x_n+\frac{2}{x_n})\]

myininaya (myininaya):

try to avoid a number such that f' =0

myininaya (myininaya):

or dont try do!

myininaya (myininaya):

also if newton's method doesn't seem to be converging than it is failing!

OpenStudy (anonymous):

yes it is nice if you can see a picture of the function you want the root of. then you can see if the approximations converge or not

OpenStudy (anonymous):

like the apple on his head. oh wait that was falling not failing. never mind

OpenStudy (anonymous):

those are synonymous :]

myininaya (myininaya):

aww thanks satellite :)

OpenStudy (anonymous):

hey i tried to tease you in a post about amistre. did you see it?

myininaya (myininaya):

satellite, while you are gone joe fills in for you

myininaya (myininaya):

no i missed it but i will looks at it

OpenStudy (anonymous):

please do. i was amused with myself

OpenStudy (anonymous):

thanks again for your help, guys. I'd be hopelessly lost without you.

OpenStudy (anonymous):

yw